OpenStudy (anonymous):

Determine an equation, in standard form, for the line that passes through (3,4) (-1,5).

OpenStudy (nenadmatematika):

y-4=(-1/4)(x-3) so (-1/4)x-y+19/4=0 is the standard form of the equation

OpenStudy (anonymous):

See I am confused because the answer says: x+4y=19

OpenStudy (nenadmatematika):

multiply equation by 4 and you'll get that answer

OpenStudy (anonymous):

Ok Thank you. I see so you just get rid of the fraction.

OpenStudy (nenadmatematika):

that's right! you're welcome :)